# EXTENSIONS OF UNIRATIONAL GROUPS

### v2

@@ 5 Deducing unirationality from quotients
We first treat the special case in which the $X_{i}$ are all open subschemes of $\mathbf{P}^{1}_{K}$, in which case we prove a somewhat stronger result – namely, that the number of maps $f$ as in the proposition is bounded in terms only of $U$ and the degrees of the complementary divisors $\mathbf{P}^{1}\backslash X_{i}$. As in the proof of Lemma 5.4 above, we may invoke [Ros1, Th. 6.7] to conclude that it is sufficient to show that, for any reduced divisors $D_{1},\dots,D_{r}$ on $\mathbf{P}^{1}$, the number of multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ is finite and bounded in terms only of $U$ and the degrees of the $D_{i}$, and this is the content of Lemma 5.4(ii).
Now consider the general case. For technical reasons, we shall find it more convenient to prove a slightly modified version of the proposition. We invoke certain notation from the beginning of [Ros3, §7]. Namely, we have associated to $f$ a map $\Delta_{f,x_{1},\dots,x_{r}}\colon\prod_{i=1}^{r}X_{i}\rightarrow U,$ which enjoys two key properties: (1) $\Delta_{f,x_{1},\dots,x_{r}}=1$ whenever any coordinate is set to $x_{i}$, and (2) if $f$ vanishes whenever any coordinate is set to $x_{i}$, then $\Delta_{f,x_{1},\dots,x_{r}}=f$. Further, the construction of $\Delta_{f,x_{1},\dots,x_{r}}$ is functorial in the obvious sense.
@@
Let $K$ be a field of degree of imperfection $r$, and let $U$ be a smooth connected unipotent $K$-group. Then $U$ is unirational if and only if $U/\mathscr{D}_{r}U$ is.
@@
Because $\mathscr{D}_{\infty}(U)=1$, we may assume that $r<\infty$. The only if direction is clear, so we prove the converse. If $U_{w}$ denotes the maximal wound quotient of $U$ (that is, the quotient of $U$ by its maximal split $K$-subgroup), then $U$ is unirational if and only if $U_{w}$ is, so we may assume that $U$ is wound. We may of course also assume that $U\neq 1$. Thus $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}$. Consider the $r$-fold commutator map $U^{r}\rightarrow U$, $(u_{1},\dots,u_{r})\mapsto[u_{1},[u_{2},\dots,u_{r}]\dots]$. This descends to a map $c\colon\overline{U}^{r}\rightarrow U$, where $\overline{U}:=U/U^{\prime}$. Because $\mathscr{D}_{r}U\twoheadrightarrow\overline{U}/\mathscr{D}_{r}(\overline{U}), U¯\overline{U}$ is unirational by dimension induction. If $c$ is constant, then $\mathscr{D}_{r}(U)=1$ and the theorem is trivially true. Otherwise, Proposition 5.6 implies that $c$ generates a nontrivial (normal) permawound $K$-subgroup $V$ of $U$. We then finish by combining dimension induction applied to $U/V$ and Proposition 5.3. ∎
@@
If $U$ is a wound unirational $K$-group, and $K$ has degree of imperfection $r$, then $\mathscr{D}_{r+1}(U)=1$ [Ros3, Th. 1.5]. (Note that the indexing on the lower central series used in that paper is unfortunately nonstandard, and in particular differs from that used here by $1$.) Therefore the quotient $U/\mathscr{D}_{r}(U)$ is in that sense not too far from $U$.
Of particular note in Theorem 5.7 is the case $r=2$, which says that if $K$ has degree of imperfection $2$, then $U$ is unirational if $U^{\operatorname{\rm{ab}}}$ is. We shall now give an example to show that this optimal, by constructing over every field of degree of imperfection $\geq 3$ a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not.
@@
We claim that $b$ generates $W$. That is, $b$ does not land in any proper $K$-subgroup scheme of $W$, or equivalently, for some $n>0$ the map $b^{n}\colon(V_{1,\lambda}\times V_{1,\mu})^{n}\rightarrow W$ obtained by adding $b$ over the $n$ coordinates is surjective. In fact, since all groups and maps live over the subfield $\mathbf{F}_{p}(\lambda,\mu)$, it suffices to verify this over this subfield, where it is [Ros3, Prop. 9.9]. (The multi-additive map given in the proof of that proposition is in this setting exactly $b$.)

### v3

@@ 5 Deducing unirationality from quotients
We first treat the special case in which the $X_{i}$ are all open subschemes of $\mathbf{P}^{1}_{K}$, in which case we prove a somewhat stronger result – namely, that the number of maps $f$ as in the proposition is bounded in terms only of $U$ and the degrees of the complementary divisors $\mathbf{P}^{1}\backslash X_{i}$. As in the proof of Lemma 5.4 above, we may invoke [Ros2, Th. 6.7] to conclude that it is sufficient to show that, for any reduced divisors $D_{1},\dots,D_{r}$ on $\mathbf{P}^{1}$, the number of multi-additive maps $\prod_{i=1}^{r}\operatorname{R}_{D_{i}/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ is finite and bounded in terms only of $U$ and the degrees of the $D_{i}$, and this is the content of Lemma 5.4(ii).
Now consider the general case. For technical reasons, we shall find it more convenient to prove a slightly modified version of the proposition. We invoke certain notation from the beginning of [Ros4, §7]. Namely, we have associated to $f$ a map $\Delta_{f,x_{1},\dots,x_{r}}\colon\prod_{i=1}^{r}X_{i}\rightarrow U,$ which enjoys two key properties: (1) $\Delta_{f,x_{1},\dots,x_{r}}=1$ whenever any coordinate is set to $x_{i}$, and (2) if $f$ vanishes whenever any coordinate is set to $x_{i}$, then $\Delta_{f,x_{1},\dots,x_{r}}=f$. Further, the construction of $\Delta_{f,x_{1},\dots,x_{r}}$ is functorial in the obvious sense.
@@
Let $K$ be a field of degree of imperfection $r>1$. Then a smooth connected unipotent $K$-group $U$ is unirational if and only if $U/\mathscr{D}_{r}U$ is.
For the corresponding result when $r=1$, see Corollary 4.5.
@@
Because $\mathscr{D}_{\infty}(U)=1$, we may assume that $r<\infty$. The only if direction is clear, so we prove the converse. If $U_{w}$ denotes the maximal wound quotient of $U$ (that is, the quotient of $U$ by its maximal split $K$-subgroup), then $U$ is unirational if and only if $U_{w}$ is, so we may assume that $U$ is wound. We may of course also assume that $U\neq 1$. Thus $U$ contains a nontrivial smooth connected central $K$-subgroup $U^{\prime}$. Consider the $r$-fold commutator map $U^{r}\rightarrow U$, $(u_{1},\dots,u_{r})\mapsto[u_{1},[u_{2},\dots,u_{r}]\dots]$. This descends to a map $c\colon\overline{U}^{r}\rightarrow U$, where $\overline{U}:=U/U^{\prime}$. (Here we use $r>1$.) Because $\mathscr{D}_{r}U\twoheadrightarrow\overline{U}/\mathscr{D}_{r}(\overline{U}), U¯\overline{U}$ is unirational by dimension induction. If $c$ is constant, then $\mathscr{D}_{r}(U)=1$ and the theorem is trivially true. Otherwise, Proposition 5.6 implies that $c$ generates a nontrivial (normal) permawound $K$-subgroup $V$ of $U$. We then finish by combining dimension induction applied to $U/V$ and Proposition 5.3. ∎
@@
If $U$ is a wound unirational $K$-group, and $K$ has degree of imperfection $r$, then $\mathscr{D}_{r+1}(U)=1$ [Ros4, Th. 1.5]. (Note that the indexing on the lower central series used in that paper is unfortunately nonstandard, and in particular differs from that used here by $1$.) Therefore the quotient $U/\mathscr{D}_{r}(U)$ is in that sense not too far from $U$.
Of particular note in Theorem 5.7 is the case $r=2$, which says that if $K$ has degree of imperfection $2$, then $U$ is unirational if $U^{\operatorname{\rm{ab}}}$ is. The same holds for $r\leq 2$, by Corollary 4.5 and because every connected linear algebraic group is unirational when $r=0$. We shall now give an example to show that this is optimal, by constructing over every field of degree of imperfection $\geq 3$ a wound unipotent group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational.
@@
We claim that $b$ generates $W$. That is, $b$ does not land in any proper $K$-subgroup scheme of $W$, or equivalently, for some $n>0$ the map $b^{n}\colon(V_{1,\lambda}\times V_{1,\mu})^{n}\rightarrow W$ obtained by adding $b$ over the $n$ coordinates is surjective. In fact, since all groups and maps live over the subfield $\mathbf{F}_{p}(\lambda,\mu)$, it suffices to verify this over this subfield, where it is [Ros4, Prop. 9.9]. (The multi-additive map given in the proof of that proposition is in this setting exactly $b$.)